40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog, Page 1 of 37  Not logged in ELOG logo
ID Date Author Type Category Subject
  1927   Tue Oct 19 13:52:03 2021 RadhikaDailyProgress1418 nm AUX ECDL1418nm ECDL Frequency noise

Attachment 1 is a diagram of the current setup for measuring ECDL frequency noise. Since the last update, I have fed the beat note signal to a mixer, using a 35 MHz LO sourced from the Marconi. The resulting demodulated signal is passed to a low-pass filter, removing the 2f sinusoidal term (any trace of the frequency difference) and leaving behind a low-frequency term containing frequency noise information of the original beam (accumulated over the length of delay line).

I took spectra of the resulting signal using the SR785 (Attachment 2). Note that these units are still in V/rtHz, since the signal has not been calibrated to the appropriate units for frequency noise, Hz/rtHz. Finding the calibration term will involve study of delay line frequency discrimination. 

Attachment 1: ECDL_diagram.pdf
ECDL_diagram.pdf
Attachment 2: ECDL_FNM_13-10-2021_151524.pdf
ECDL_FNM_13-10-2021_151524.pdf
  1926   Mon Oct 4 17:44:34 2021 RadhikaDailyProgress1418 nm AUX ECDLFree space AOM

[Paco, Radhika]

Last Friday we received a new lens to direct the AOM 1st-order beam from free space into a fiber cable. We mounted the lens and connected a fiber cable into the photodiode, and tried to align the lens and see a jump in the oscilloscope. We were not able to do so and wrapped up for the day.

Today we continued aligning the lens with the fine adjustment on the mount, and eventually saw signal on the scope! Hooray, done with free space. We then prepared for eventually taking a heterodyne beat note measurement and hooked up the appropriate inputs/outputs to the beamsplitters. We added in the 50-50 beamsplitter that takes in the 1st order diffracted beam along with the beam from the delay line as inputs. We passed one of the outputs to the photodiode and had to retweak the freespace-to-fiber lens until we recovered signal on the scope, and we saw the beatnote signal.

Next, while Paco is out of town I will continue to work towards making a frequency noise measurement. We made a roadmap today:

I will demodulate the beat note using a mixer and a 35 MHz LO sourced from the Marconi. The result will be a 2f cosine term, along with a much lower frequency term which encloses the frequency noise information. This will be passed through a low-pass filter to get rid of the first high-frequency term. The remaining time-domain signal will be passed to the SR785 to obtain a spectra of the frequency noise. Calibration will need to be performed to obtain the right units for the spectra, Hz2/Hz (or Hz/rtHz). 

  1925   Wed Sep 22 16:44:34 2021 RadhikaDailyProgress1418 nm AUX ECDLFree space AOM

[Paco, Radhika]

We had previously noticed that the ECDL laser power seemed weaker compared to when we originally set it up and tested it. Today Paco opened it up and tweaked the grating inside to obtain a max power of 3 mW. This way, we could better resolve the 0th and 1st order beams coming out of the AOM.

Since we don't yet have a lens to send the collimated 1st-order beam to fiber, we connected a power meter to detect the beam and hooked it up to the oscilloscope. We noted peak diffraction at around 38.5 MHz (rough estimate). Using the inverse relationship between laser wavelength and the RF frequency f \lambda = constant, and the fact that the AOM is designed to operate at 1550 nm at 35 MHz, we calculated that the ECDL wavelength should be ~1409 nm. Of course this is a rough estimation, but it is a quick validation that we are indeed operating near 1418 nm. 

  1924   Thu Sep 16 15:21:21 2021 PacoDailyProgress1418 nm AUX ECDLFree space AOM

[Paco,  Radhika]

Uninstalled the fiber AOM and temporarily removed the third fiber 2x2 port beamsplitter. We are now using this free-space AOM. Then, I managed to launch one of the outputs of the second fiber beamsplitter into free space using a F220APC-1550 fixed collimator. The beam clears the  AOM aperture nicely and lands on the other side.

This AOM operates at a RF frequency of 35 MHz, so we set up a sweep on the Marconi to cover a window of 35 MHz +- 15 MHz. Using an IR detector card, we looked for evidence of 1st-order diffraction (from the setup geometry, the 1st order beam should have been visibly discernable). We first scanned the AOM across yaw but did not notice diffraction. Then, Paco lowered the height of the fixed collimator and we repeated scanning across yaw. We eventually saw the beam "jump" - diffraction! We adjusted yaw until we recovered both 0th and 1st order beams, at 50/50 intensity.

In summary, the free-space AOM works and we have managed to see 1st order diffraction. Next steps will be to quantitatively measure this diffraction while sweeping across RF frequency and power.

  1923   Thu Sep 2 17:31:38 2021 RadhikaDailyProgress1418 nm AUX ECDL1418 nm ECDL Relative Intensity Noise

I took a relative intensity noise (RIN) measurement of the ECDL, by feeding the 0th order output of the AOM to the SR785. The RF power driving the AOM was set to 0 dBm. The RIN at 1 Hz is about 3x10-5, which is consistent with informal measurements we took on 08/13. From my understanding this noise looks pretty low, which is good. I will consult with Paco and add more discussion or conclusions, if any.

Attachment 1: ECDL_RIN_02-09-2021_165151_alone.pdf
ECDL_RIN_02-09-2021_165151_alone.pdf
  1922   Wed Sep 1 13:12:02 2021 RadhikaDailyProgress1418 nm AUX ECDL1419 nm ECDL AOM diffraction at 95 MHz

[Paco, Radhika]

Today we tried to pick up from [1920] by repeating the sweep measurements across RF frequency, at 3 dBm (max power). We noticed that the 0th order signal would dip around the expected value, consistent with the plot in [1920]. However, there was no signal from the 1st order. Clearly diffraction was occurring as seen by the dip in 0th order, but nothing was coming out of the 1st order port. We spent some time debugging by swapping the photodetector inputs / playing with the PD gains / performing power cycles, but got no insight into the issue. 

We suspected the 1st order fiber coming out of the AOM might be damaged, since it loops around fairly tightly. After giving it more slack, we still saw no signal. We wanted to test the fiber, so we took an unused output of the 50-50 beamsplitter and fed it into the 1st order port, effectively running the AOM in reverse. We hooked up the input and 0th order ports to the photodiodes and did not observe any signal. From here we were more convinced that the 1st order fiber may have seen some damage. 

For next steps, we can still use the existing fiber setup to take measurements of relative intensity noise (RIN), using the 0th order output of the AOM. I plan to do this in the next few days. Meanwhile, Paco is looking into ordering parts for a free space setup. We found a free-space AOM at 1064nm that seems promising, and we will work to transition the setup accordingly. 

  1921   Tue Aug 17 11:09:29 2021 ranaDailyProgress1418 nm AUX ECDL1419 nm ECDL AOM diffraction at 95 MHz

Should measure the S-matrix using a bi-directional coupler.

  1920   Thu Aug 12 11:49:59 2021 RadhikaDailyProgress1418 nm AUX ECDL1419 nm ECDL AOM diffraction at 95 MHz

[Paco, Radhika]

When previously trying to characterize the AOM, we had noticed no 1st order diffraction when operating at 80 MHz, but significant diffraction at 95 MHz. This motivated us to take measurements while sweeping across both RF drive frequency and Marconi drive power. For frequency, we swept from 80-120 MHz in steps of 1 MHz. For power, we swept across [3, 0, -3] dBm (3 dBm is max power before saturating AOM). We took our measurements of 0th and 1st order signal using an oscilloscope.

Contour plots of the 0th and 1st order signals can be seen in Attachments 1 and 2, respectively. Peak 1st order diffraction seems to occur at ~106 MHz. Using this AOM for a beat note measurement, the frequency difference would be greater than intended, which could lead to a weaker beat note signal. 

*Bonus: Today we moved the ECDL setup off the cryostat table and onto the other table. These measurements were taken after the move. 

Attachment 1: zeroth_order_contour.pdf
zeroth_order_contour.pdf
Attachment 2: first_order_contour.pdf
first_order_contour.pdf
  1919   Tue Aug 10 11:00:43 2021 PacoDailyProgressDOPODOPO v2

[paco, nina]

We started rebuilding the DOPO in the lab even though the new optical table hasn't arrived. For this reason, we are using a 1 ft x 3 ft x 0.5 in anodized aluminum breadboard which we can then attach handles to move the setup. This also makes the prototype's footprint smaller. The first thing we did as usual was settle on a beam height. The beam height used before was ~ 3in (~ 75mm), and since the EOM, Faraday Isolator, and RFPD are nominally at that height from the breadboard, the only thing we had to fix was the pump laser head. The bare height is 55 mm, so we stacked two 9 mm thorlabs bases together, bolted them down to the breadboard and then mounted the NPRO laser head on the top. Finally, using a level we secured it to the breadboard using the three points and long 1/4-20 screws while being careful as we didn't want to flex the head too much.

Next up is aligning the laser to the EOM and Faraday Isolator. For this, we will use the beam profiles measured late last year. Another task ahead is to implement the new mount for the cavity.

  1918   Tue Aug 10 09:51:44 2021 PacoDailyProgress1418 nm AUX ECDL1419 nm ECDL AOM diffraction at 95 MHz

[radhika, paco]

We changed the setup to use a low power amplifier rather than the 5W amp from last time. The updated schematic is in Attachment 2. This is in part because 5W is an overkill to drive a fiber AOM which is known to saturate at 0.6 mW of RF input, but also because working with lower power active elements is easier and considerably safer. We dropped the 5W amp. in Rana's office last Friday, and got a ZHL-3A-sma. This little guy gives a max power output of 29.5 dBm (~890 mW) which should be more than enough while using the Marconi as our source (max output +13 dBm).

We hooked the amplifier to the load (AOM) without any couplers or attenuators in between, powered it with +24 VDC and quickly repeated a scan of the source power level while to see any sign of diffraction in the PDs. The result is in Attachment 2. We were a little bit disappointed that there appeared to be no diffraction, so next we tried scanning the RF frequency (it was nominally at 80 MHz) around and we finally succeeded in seeing some diffraction at 95 MHz! Paco thinks the internal fiber coupling made for the design wavelength (2004 nm) is suboptimal at 80 MHz and ~1.4 um wavelength. Therefore, to couple the 1st order back into the fiber, we need to shift the RF frequency to restore the diffraction angle at the cost of potentially not driving the optimal efficiency. An interesting observation made at the same time we saw 1st order light was that the power seemed to drift very slowly (-1%/min), which may have to do with some thermal drift inside the crystal... Our plan is to make a complete characterization of the diffraction efficiency at 1.4 um, and also investigate the slow intensity drifts as a function of RF input. The goal is not so much to understand and fix this last one, but to be able to operate the setup at a point where things are stable for a low frequency, frequency noise measurement.

Attachment 1: rf_setup.jpg
rf_setup.jpg
  1917   Wed Aug 4 11:36:30 2021 RadhikaDailyProgress1418 nm AUX ECDL1419 nm ECDL with 2um AOM tests

[Radhika, Paco]

In order to transition the ECDL laser noise characterization to a heterodyne setup, we needed to test the AOM (acousto-optic modulator). We wanted to drive the AOM at 80MHz using the Marconi signal generator. Since the AOM has a max driving power of 600 mW, we determined that if we run the Marconi at max output power (13dBm), we saturate the AOM through a variable attenuator and a 5W amplifier. The detailed setup is in Attachment 1.

As we scanned the AOM RF input power, we monitored the mean of the 0th and 1st order power outputs using 2 amplified photodiodes on the scope. Attachment 2 plots the results of the scan; although we noticed the 0th order dropping, we did not see evidence of diffraction in the 1st order. Our suspected theory is that the lost power from the 0th order is due to thermally-driven attenuation inside the AOM (we do not know what is inside the AOM, so this is purely speculative). The next thing we want to try is to add a DC power level to the AOM RF input, but we will double check with Aidan.

Attachment 1: rf_setup.jpg
rf_setup.jpg
Attachment 2: diffraction_levels.png
diffraction_levels.png
  1916   Wed Jul 14 12:21:08 2021 PacoDailyProgress1418 nm AUX ECDLECDL lases... and MZ locked mid-fringe

[Paco]

Worked for a few hours to get the aspheric properly aligned. The procedure is quite finnicky, as the four 2-56 flexure screws have too much game and the fine thread setscrew that adds tension is too constrained. Anyways, it generally goes like this:

  1. With the grating out of the way, and the 2-56 screws slightly loose, move the aspheric flexure until a collimated beam (as round as possible) exits the centered round aperture in front of the SAF chip.
  2. Very carefully tighten the fine threaded setscrew in place to register the aspheric alignment.
  3. Check that the desired beam hasn't changed
  4. Insert grating careful not to touch the aspheric flexure (again, the mechanical registration is not great!)
  5. Manually rotate the grating while monitoring the power at the fiber output until non-ASE light appears. This is quite sensitive to alignment/angle and the better the mode is matched back into the fiber, the easier it is to find the right position.
  6. Fix the grating flexure.
  7. Slightly tweak the grating mirror knobs by hand to maximize said power (careful to avoid saturating PD)

After this, I installed a second amplified InGaAs detector, hooked up the unbalanced MZ beamsplitter output into the two PDs, adjusted the gains to equalize the output voltages and then hooked the two signals to the A and B inputs of an SR560 in "A-B" mode. The output (gain 1) was good enough to feed back in the HV PZT amplifier input modulation which allowed the MZ to lock mid-fringe. The lock is rough, as the balanced homodyne signal retains a tiny offset due to imperfect balancing... Attachment 1 shows the setup, including a typical scope trace after coarse current tuning (Ch1 and Ch2 in yellow and blue represent the photocurrents in the two MZ ports in the absence of feedback).

Indeed, scanning the nominal PZT voltage broke the lock, potentially after crossing a mode hopping region.

Tasks to be done:

  • Power characterization, including RIN, emission vs current.
  • Emission spectrum characterization; using the homemade spectrograph (grating + lens + camera) as the PZT is scanned
  • MZ feedback loop characterization, including optimizing the balancing stage
  • Self-homodyne phase noise spectrum

Next, as was suggested during yesterday's group meeting, we will transition into a self-heterodyne setup (with an AOM which I have yet to check out in the QIL).

Attachment 1: selfhomodyne_1418.jpg
selfhomodyne_1418.jpg
  1915   Fri Jul 9 11:40:42 2021 PacoDailyProgress1418 nm AUX ECDLNew aspheric flexures

[Radhika, Paco]

We tweaked the flexure alignment until we had a nominally collimated beam (~2 mW @ 250 mA of diode current) through the output aperture in the ECDL housing. We noted that the collimated beam is off-centered on that circular aperture along the horizontal (yaw) angle. After this, Radhika installed the ECDL grating and we hooked up the fiber output onto a InGaAs PD to monitor the power output. We tweaked the alignment of the grating (mostly yaw) to try and see a change in the power output to indicate optical gain in the diode, but saw no changes. We observed a change in the PD photocurrent as a function of the diode current in the absence of the grating (no optical feedback) which is indicative of ASE. We measured this level to be ~ 140 mV at 200 mA of current; with no observed threshold. In conclusion, we still need to refine our grating alignment to provide gain on the diode and observe lasing at the nominal 1450 nm wavelength.

Attachment 1: IMG_1773.jpeg
IMG_1773.jpeg
  1914   Wed Jul 7 13:05:07 2021 PacoDailyProgress1418 nm AUX ECDLNew aspheric flexures

[Radhika, Paco]

Today we fired up the 1418 nm ECDL and attempted initial adjustment of the aspheric lens. The design follows D2100115 which is a copy of the 2 um ECDL so we just changed the diode, the grating flexure angle, and the aspheric + flexure assembly and we are good to go. Radhika removed the 1900 nm aspheric flexure and we mounted the new collimating assembly which uses a f=3.1 mm (NA = 0.69) lens. At the beginning we had to feed over 300 mA of current to be able to see a beam (which was still diverging) so we had to free the flexure completely and align by hand to find the nominal positioning for a collimated beam. We lost a 2-56 screw in the process, but the final assembly is still in progress. The plan to follow is:

  • Finalize flexure alignment
  • Insert grating
  • Characterize ECDL emission
  1913   Thu Jun 10 09:59:52 2021 PacoLab InfrastructureDOPODisassembly for new optical table

Today the DOPO v0 got disassembled to make way for the optical table swap. Most components have been stored in the white cabinet's bottom panel.

Attachment 1: IMG_20210610_092713.jpg
IMG_20210610_092713.jpg
  1912   Wed Jun 2 18:37:09 2021 PacoSummaryCrackleVent crackle experiment

[Paco, Anchal, Ian, Yehonathan]

Today, in preparation for the optical table to come, we vented the big crackle jar using the vent valve near the gauge. We detached the roughing pump and covered the bellows and pump connections with clean aluminum foil. We then proceeded to move several instruments, including some other pumps, a compressor, a couple of power supplies, power cords, the HeNe laser, misc. material blocks, and boxes with bearings and springs into the cage. The next operation required for us to displace the table is to lift the jar from the top and carefully dismantle the Crackle experiment and store it away somewhere.

Questions: where to store mostly?

  1911   Thu May 20 17:09:43 2021 KojiSummaryGeneralAnother Heimann Sensor

Another Heimann Sensor / Boston Electronics delivered to Paco.
This unit (purchased May 2020/ / Delivered Aug 5th, 2020) has a FZ-Si window on it.
We don't know how it is.

Attachment 1: P_20210520_151126.jpg
P_20210520_151126.jpg
  1910   Wed May 19 09:25:34 2021 PacoNoise HuntingDOPO316 Hz noise

[Paco]

- Have been investigating 316 Hz noise in the control signal for the DOPO lock. Here is a list of some things that have been ruled out, mostly electrical:

  - EOM power supply --> noise still present in DOPO transmission
  - RFPD DC out --> no funky ground loops with scope (also looking at demod signal in different channel), noise still visible in transmission
  - RFPD power supply --> noise still visible in transmission...
  - Pump laser intensity (upstream pickoff) --> not a great test because pickoff optics are also on the optical table..
  - 2 x SR560s --> No effect after bypassing
  - Marconi --> same result as with anything in the loop after RFPD demod

- Things left to rule out:
  
  - Fume hood exhaust fan ** highly suspected, my phone's own cheap-o microphone power spectrum shows peaks at 316.5 Hz (!) when near the exhaust fan
  - NPRO temp controller fan --> phone audio spectrum shows line noise (60 Hz) mostly, and also 188 Hz... need to test further independently of the fume hood...

In ruling out the 6-axis translation mount on the DOPO cavity, I removed the PPKTP crystal + oven temporarily but still saw the noise. Since the resonator was no longer stable without the crystal, I needed to bring the mirrors closer and realign the output coupler from scratch. 

Restored DOPO cavity with crystal, alignment. MM efficiency ~ 35%... still optimizable.

  1909   Tue May 18 10:28:50 2021 PacoLab InfrastructureEquipmentLoanheimann sensor update

Heimann (HTPA80x64d) thermopile array;

- First test to grab frames was done in my personal Win10 machine, with no success. Either I was unable to configure the server correctly, or the software "ArraySoft" is not supported in Win10. Upon contacting Heimann, I received instructions to update to a newer version but was warned that it's just a new GUI, nothing really changed from v1 --> v2. So didn't even bother.

- Instead, wrote a simple python-socket UDP server to catch the video stream. Most trouble happened when using temperature mode (command "K"). The client streams a bunch of zeros... My guess is that this unit does not have an internal temperature calibration, and one could in principle be uploaded but we probably don't care. Streaming in raw voltage mode (command "t") works well, as shown by the sample frame shown in Attachment 1.

- After recovering the CTN Win7 laptop from Radhika, I gave "ArraySoft" another change just to know the frames I was getting in python were not bogus. For this I pointed a 532 nm laser pointer straight to the sensor and got an image shown in Attachment 2. The key difference is the processing of the video stream. Attachment 1 is a single frame, while Attachment 2 is the average of 30 frames with no offsets present. 

- Another issue present during this task was a faulty USB connection. Sometimes moving the sensor around would interrupt the stream (power lost). I carefully removed the case and exposed the TO-39 package and surrounding electronics to inspect and search possible failures but after seeing none, I swaped the USB power cable with my portable battery charger and had a more robust operation... So I dumped the old USB cable, and will get a new one.

- Since this one was borrowed from TCS lab, I placed an order for another one which will be set up permanently in the lab. Hopefully this will be enough for the OSA.

Attachment 1: no_light.png
no_light.png
Attachment 2: light.png
light.png
  1908   Wed Apr 14 16:49:30 2021 PacoMiscOptomechanicsDOPO mount v2

Drew some new mounting scheme for the DOPO cavity; main revisions with respect to the current mount are -->

  • Side mounts for both mirrors (instead of vertical)
  • Both mirror mounts are the same (3-axis polaris K1) so both mirrors need to be attached properly
  • Improved access to align crystal using newport 9031 (6 axis displacement mount), which is crucial to make the DOPO fields co-resonant

Attachment 1 illustrates the design; shows three views of the same assembly.

Concerns: mechanical noise from side mounted mirrors ... for this, there could be a solid piece which makes a rigid connection between the two mirrors (that's why they are upside down) and perhaps between the two tall posts (so S-shaped as viewed from the top)? Still working on this.

Attachment 1: dopo_mount_v2.pdf
dopo_mount_v2.pdf dopo_mount_v2.pdf dopo_mount_v2.pdf
  1907   Tue Mar 30 15:18:06 2021 PacoLab InfrastructureEquipmentLoan2um CCD borrowed

With Aidan's assistance, I borrowed

  • WiDy SWIR camera (Pembroke WiDy SWIR 640U-A) from the QIL (Attachment 1)
  • Heimann thermopile sensor (HTPA80x64d) with microUSB-A and ethernet cables from the Adaptive optics lab (Attachment 2)

for ~ 2 um imaging in the Crackle lab.

Attachment 1: IMG_20210330_151046.jpg
IMG_20210330_151046.jpg
Attachment 2: IMG_20210330_150958.jpg
IMG_20210330_150958.jpg
  1906   Tue Mar 9 19:21:38 2021 PacoMiscOpticsDOPO cavity pole

- Noticed that the cavity transmission peaks @ 1064 nm were much wider than originally estimated by the dopo cavity design notebook suggesting a lower Finesse. So using the PDH error signal, and knowing the EOM sidebands are at 36 MHz estimated the current DOPO cavity linewidth to be 19.5 MHz, well in excess of the target 10.4 MHz.

- Updated the crystal AR coating specs from Raicol (R < 0.3% @ 1064/2128), but more importantly, I included the absorption coefficient of KTP, alpha=0.005/cm (often quoted as < 0.01 / cm) into the roundtrip loss and the design now gives 17.97 MHz. So, given the uncertainty in the absorption coefficient of the NL crystal, and all the coatings in the experiment, this adjustment might be enough to explain this observation.

  1905   Thu Feb 25 10:28:07 2021 PacoDailyProgressOpticsDOPO lock endurance

Test long term stability of the DOPO cavity lock; The cavity remained resonant overnight (start ~ 8 PM yesterday) and lost around 11 AM today. It might be good enough to approach lock point manually using laser temp. control and then engage the fast loop. In any case, today will set up an acromag channel for this. Configured "XT1541-2um-SlowDAC" to 10.0.1.47

  1904   Wed Feb 24 17:42:49 2021 PacoDailyProgressGeneralDOPO locking and SHG

Locking update

The plan during these past few days has been to have fast control loop of the cavity (locked to laser using PZT, which succeeded using SR560s), and slow control loop where the laser temp. actuator is fed back the integrated PZT input to follow the long term cavity drift. For that, have been messing around with the high-level (GUI) API of PyRPL, with basically no success. In fact currently the RedPitaya cannot even replace the SR560 fast controls, which probably has to do with the +- 1 Volt limits on the RP input/output.

Another issue is that any loop gain depends on the REFL power, which will be at some point slowly ramped up to cross the OPO operating threshold, and while there is a (PBS + HWP) knob on how much light is hitting the RFPD, the lock is not yet good enough to keep up with the slow human action.

First light from nonlinear conversion

WIth the cavity locked, and under ~ 220 mW of pump (right before the cavity, i.e. 1.3 Amps of current on the driver), noticed a tiny green dot coming from within the crystal oven. This is pretty great news in terms of phase matching, but not necessarily so in terms of the right parametric conversion process (understanding is that SHG is easier to attain even with single pass). See tiny green spot as caught using phone camera in the attachment.

Attachment 1: shg.jpg
shg.jpg
  1903   Thu Feb 18 09:39:01 2021 PacoDailyProgressElectronicsPDH error signal misbehaving

Error signal

Upon closer inspection the error signal seems to vary quite significantly on the scope (scanning @ 2 Hz), sometimes completely flipping its sign even though it always triggers on the same side of the ramp (see attachment for video, along with some neck excersise).

This might be the same behaviour from before, whereby the demodulated signal might still be "riding" a low-freq componen which can't be compensated with the LO (Marconi's carrier resolution = 1 Hz). Using the 10 MHz external Rb reference doesn't change anything. It seems that even with the coupler, reflections may be entering the mixer... 

Adding a LP filter (BLP-1.9+) right at the mixer output solves this for good. Even using 36 MHz LO vs anything else doesn't make a difference so this explains the previous issue. Moving back to lock using stable err signal.

For reference, the LO carrier is set to 36.000 MHz, +7 dBm (so the EOM is driven with an estimated +30 dBm well below the saturation or damage threshold +40 dBm).

DOPO locked

Achieved a good lock for pretty much all of the afternoon today. The laser ran at 937 mA current, the optimal gain on SR560 was found to be 50, with a LP cutoff at 300 Hz (12 dB/oct rolldown). The 300 Hz cutoff supresses most of the nasty 8 kHz noise (and harmonics) which I can hear with enough gain. Source still to be determined.

Attachment 1: misbehaving_pdh.mp4
  1902   Wed Feb 17 11:56:48 2021 PacoLab InfrastructureElectronicsUPDH box zero model and SR560 "lock"

UPDHv3 box (serial 17142) is bogus. While retrieving values of some of the components to plug into working zero model, saw the VGA stage is bypassed by a previously unnoticed hack. Verified this by taking TF and not seeing any changes with respect to the gain knob (shown below are zero's model TFs suggesting a tunable UGF from ~ 10 Hz to 1 kHz), so this box is not good for a standalone servo.

As suggested a few meetings ago, made a quick and dirty lock using a single SR560 and took measurement of something* CLTF (SR560 gain = 10) below. New goal is to find a decent replacement, for which decided to use RedPitaya's python API "pyRPL". Just using the GUI out of the box can also lock the cavity relatively quickly but neither method results in longer than 1 minute lock... so took one step back to polish the pdh error signal.

* Something = Use SR785 TF measurement with source on Ch1, and to B input in SR560. The SR560 in (A-B) mode, and demodulated signal connected to A. The loop was closed with the SR560 output driving the PZT, and Ch2 of SR785. Wouldn't call this CLTF...

Attachment 1: updhv3_VGA_gain.pdf
updhv3_VGA_gain.pdf
Attachment 2: SR560_OLTFSR785_17-02-2021_164500.pdf
SR560_OLTFSR785_17-02-2021_164500.pdf
  1901   Fri Feb 5 14:15:22 2021 ranaDailyProgressOpticsOPO cavity lock

For the splitting, I recommend not to use a splitter.

Instead, you can use a -10 or -20 dBm bi-directional coupler. You send the -10 dBm signal to the EOM amp, and you can fill up the needed power for the LO mixer. Also the "bi" nature of the coupler means that you can check for reflected power to diagnose if you are having impedance mis-match. Since you don't have an isolation amplifier in your setup, its important to make sure that reflections from one leg don't go back into the oscillator and disturb the other leg. Or maybe your oscillator box has an isolation amplifier between the oscillator and the splitter?

  1900   Thu Feb 4 16:46:06 2021 PacoDailyProgressOpticsOPO cavity lock

Demodulation stage

Update on demod. for OPO cavity lock. Last related elog entry described prevalence of <= -77 dBm of odd line noise harmonics (60, 180...) Hz, along with poor SNR PDH error signal. First attachment is a drawing of the current RF connections. Upon completing list of suggested actions from this post, the difference was mostly made by looking at RFPD RF out power before mixer < -40 dBm. This was no good, so after realizing that the OD = 3 nd filter before RFPD was only allowing 80 uW of a nominally reflected ~25 mW, swapped the ND filter with HWP + PBS for adjustable power splitting. Then, a healthier  -10dBm made it into the mixer and SNR improved considerably (see second attachment). Upon closer examination of err signal, low freq. sinusoidal modulation sat on top of it suggesting slightly off-resonant demodulation so finely adjusted the (Marconi) LO frequency from 36.000 MHz --> 35.999828 MHz until the error signal had a good enough shape (see third attachment below).

Lock

First attempt at cavity lock was done with ~46% mode matching efficiency and max. modulation depth (estimated ~0.21) on the EOM. The loop is achieved using UPDH box (v3) which I stole from CTN lab. Upon connecting all the inputs, scanning the phase shifter without making much of a difference, and enabling the lock, saw a stabler higher order mode on the cavity transmission which is nice. The natural follow up of scanning the PZT driver (i.e. as an offset) and re-engaging the lock resulted in what I can only describe as a "visit to the dentist", where the cavity PZT (on the output coupler) was resonating quite loudly (!!). After looking at the output monitor of UPDH box with engaged lock on SR785 an ~ 8 kHz peak explains the noise as an audible mechanical resonance. Adjusting the servo gain finely tunes it out a bit, and adding an SR560 in line before the PZT driver unit greatly helps, but changes the overall loop gain and the lock becomes unstable...  Current efforts are therefore geared towards improving the pdh loop, for which an option is to bypass the thorlabs MDT694 HV piezo driver and directly connect the UPDH output to PZT (which it may be meant to directly drive) and use slow temp. control on pump laser to approach the lock point. Another option, involving way more time, would be to *not* use UPDH box at all and implement a digital feedback loop + filter with the Red Pitaya. Perhaps the pragmatic action is to get the analog solution working and develop digital solution on the side.

Attachment 1: rf_diagram.png
rf_diagram.png
Attachment 2: offresonantpdherr.jpg
offresonantpdherr.jpg
Attachment 3: pdherrorsig.png
pdherrorsig.png
  1899   Tue Feb 2 17:39:52 2021 PacoDailyProgressOpticsre: Figuring out how much astigmatism is hurting us

Motivated in part by the conclusions below, improved estimated mode matching efficiency from a poor 13% at the beginning of day to 48% (estimated using the reflection signal levels from the rfpd). What helped was walking the beam with the last two mirrors, and then scanning the cavity output coupler around to center the resonant mode which at this point seems optimal. This process was tedious, but effective apparently.

The distance between the two mirrors is ~ 45 mm which slightly undershoots the planned 47.5 mm which could limit the achievable 100% in simulation-land, but I'm moving on for now, hoping the lock will bump it up enough for the OPO threshold to be within our pump power range.

Quote:
  • Astigmatism should not be hurting us significantly.
  • The mode matching in principle can be improved in the experiment
  1898   Tue Feb 2 10:32:25 2021 AnchalSummaryOpticsFiguring out how much astigmatism is hurting us

Methods

  • Use the actual measured beam profile in X and Y directions.
  • Propagate them with the current position of lens.
  • Assume the position of cavity mirror and crystal as given by the second solution in BestSolutions.mat in the Jan 8th analysis which is implemented currently.
  • Calculate the overlap with target and position of waists in X and Y direction.

Conclusions

  • Astigmatism should not be hurting us significantly.
  • The mode matching in principle can be improved in the experiment.

Analysis

Attachment 1: BeamProfilePropagation.pdf
BeamProfilePropagation.pdf
  1897   Tue Jan 26 11:47:12 2021 PacoDailyProgressGenerallow quality PDH error signal

After getting what looked like a decent cavity reflection signal, installed RFPD yesterday. For this, removed the lens that was right before the PD because the RFPD area is large enough, but keep ND filter in place. Powered with +- 18 VDC and monitor DC out on the scope, and RF out is sent to the IF of the mixer in the PDH box. For the LO, split the Marconi RF output and connected the demodulated signal into Ch2 of the scope in hopes that there was an error signal.

A hint of the error signal is present (blue trace below), although deeply buried in line noise (and harmonics up to ~180 Hz) so there really are two things to optimize now -->

  1. Line noise (hunting for ground loops or equipment, e.g. power supplies, analyze LO spectrum before/after splitters, mixers, etc...)
  2. Mode matching (this was the first reaction, as improving the cav refl SNR by means of mode matching makes a better pdh err signal)

Other things attempted so far -->

  • Switched mixers, splitter, and RF cables
  • Bypass the phase shifter completely
  • Scan LO phase
  • Floated RFPD power supply
  • Floated PDH box power supply (really only affecting the phase shifter if anything, though unlikely to matter at this point)
Attachment 1: poor_PDH_err.jpg
poor_PDH_err.jpg
  1896   Tue Jan 26 11:34:51 2021 ranaDailyProgressOpticsIncluded lens made by cavity input mirror and distrotion due to crystal

would be good if you could find a solution that is not very sensitive to precise lens placement

 

  1895   Tue Jan 26 11:33:32 2021 ranaDailyProgressOpticsIncluded lens made by cavity input mirror and distrotion due to crystal

Quote:

Error in previous calculations:

  • We did not take into account the effect of cavity input mirror on the coupled light. It would act as a thick concave lens for the coupled light into the cavity.
  • We did not take into account the divergence due to refraction at the crystal surface.

Goals and restrictions:

  • Use the fewest lenses as possible after having used a fixed lens at 67" point before the faraday isolator.
  • Choose a lens from a list of focal lengths available in west bridge labs.
  • Find the best overlap with the target beam of 18 um at the cavity waist with the most sensitivity with respect to lens positions.
  • The lens should not be closer than 1.5 inches from each other.
  • The beam widths should not exceed 4mm in diameter anywhere to ensure small areas of lenses are used.
  • Take into account the concave lens due to the input mirror.
  • Take into account the refraction due to crystal surface.

Analysis & Results

  • CavityLens.m is run to try all possible lens combinations for 1-lens or 2-lens solutions using ../20201222_BeamProfileNeatEOM/SeedBeam.mat as the seed beam.
  • The cavity input mirror is modeled as two refracting surfaces separated by 6.5mm. The first surface is flat while the second has ROC of -25 mm.
  • The crystal is modeled as two refracting flat surfaces separated by 20 mm.
  • The target beam waist is kept at the center of the crystal with 35.578 um diameter.
  • Then save all possible solutions with more than 90% overlap and where lenses are atleast 1.5" away from each other in AllPossibleSolutionsAbove90.mat using findPossibleSolutions.m.
  • findBestSolutions.m increases the overlap threshold to 0.995, allows maximum beam radius of 2mm anywhere and plots the best solutions in order of positional sensitivity of the lens. These are stored in BestSolutions.mat.

Analysis & Data

 

  1894   Fri Jan 15 17:21:53 2021 PacoDailyProgressElectronicsRFPD TF

Record TF for RFPD SN09, resonant at 36 MHz, using the exact procedure as with EOM Resonant Driver.

Attachment 1: RFPD009_TF_Inp_to_RFMon_Wide_15-01-2021_172008.pdf
RFPD009_TF_Inp_to_RFMon_Wide_15-01-2021_172008.pdf
  1893   Fri Jan 15 10:57:39 2021 PacoLab InfrastructureEquipmentLoanBorrow AG4395A and EOM Driver TF re-tune

See equipment borrowing note here.

Attempting TF measurement for resonant EOM driver, but not having luck reproducing the measurements done recently (Dec-03), so I started debugging the circuit. Both power supply connections (+- 18 VDC) seem nominal. The MAX2470 buffer regulated input is nominal at 5VDC. Looking at MMBT5551 HF transistor, base-emitter voltage is -0.60 VDC (nominal wrt -0.66 V). Using a scope, I feed a single tone (36 MHz, 190 mVpp) and look at the RFmon output and it looks ok (gain ~ 1). I changed the RFmon SMA cable and that seemed to do the trick... Bad cable (now in trash) stole my morning.

Tune EOM driver resonance to 35.993 MHz (shown below for reference).

Attachment 1: EOMDriver_TF_Inp_to_RFMon_CloseUp_15-01-2021_160402.pdf
EOMDriver_TF_Inp_to_RFMon_CloseUp_15-01-2021_160402.pdf
Attachment 2: EOMDriver_TF_Inp_to_RFMon_Wide_15-01-2021_151039.pdf
EOMDriver_TF_Inp_to_RFMon_Wide_15-01-2021_151039.pdf
  1892   Wed Jan 13 17:52:35 2021 PacoDailyProgressOpticsOPO cavity resonance

Observed first resonant transmitted (& reflected) light from the DOPO cavity; the PZT scan was centered at 31 V, at 2 Hz, with an amp. of 1.5 Vpp. To get there, revisited the path's alignment upstream to the last mirror (before the last lens), removing, inspecting, and reinstalling each component. After this, I used the camera at the end of the optical path as a "pinhole" (beam center placeholder) and after inserting each element (mirrors / crystal) checked carefully that the beam was landing straight. Then, patiently scanned various knobs (mirror mounts X/Y/XY, crystal manually) until HOM started resonating. After a bit of further alignment managed to see transmission dips in the FI pickoff. Below are two photos illustrating the current state (way more optimization is needed), as well as the setup viewed from one side (for the scope picture, purple is the ramp, yellow is cavity reflection, green is cavity transmission). Will keep optimizing in the couple next days, all at low power first, and then start cranking the power up to factor in any thermal effects into the optimized cavity.

Attachment 1: dopo_first_resonance.jpg
dopo_first_resonance.jpg
Attachment 2: dopo_sideview.jpg
dopo_sideview.jpg
  1891   Fri Jan 8 17:17:19 2021 AnchalDailyProgressOpticsIncluded lens made by cavity input mirror and distrotion due to crystal

Error in previous calculations:

  • We did not take into account the effect of cavity input mirror on the coupled light. It would act as a thick concave lens for the coupled light into the cavity.
  • We did not take into account the divergence due to refraction at the crystal surface.

Goals and restrictions:

  • Use the fewest lenses as possible after having used a fixed lens at 67" point before the faraday isolator.
  • Choose a lens from a list of focal lengths available in west bridge labs.
  • Find the best overlap with the target beam of 18 um at the cavity waist with the most sensitivity with respect to lens positions.
  • The lens should not be closer than 1.5 inches from each other.
  • The beam widths should not exceed 4mm in diameter anywhere to ensure small areas of lenses are used.
  • Take into account the concave lens due to the input mirror.
  • Take into account the refraction due to crystal surface.

Analysis & Results

  • CavityLens.m is run to try all possible lens combinations for 1-lens or 2-lens solutions using ../20201222_BeamProfileNeatEOM/SeedBeam.mat as the seed beam.
  • The cavity input mirror is modeled as two refracting surfaces separated by 6.5mm. The first surface is flat while the second has ROC of -25 mm.
  • The crystal is modeled as two refracting flat surfaces separated by 20 mm.
  • The target beam waist is kept at the center of the crystal with 35.578 um diameter.
  • Then save all possible solutions with more than 90% overlap and where lenses are atleast 1.5" away from each other in AllPossibleSolutionsAbove90.mat using findPossibleSolutions.m.
  • findBestSolutions.m increases the overlap threshold to 0.995, allows maximum beam radius of 2mm anywhere and plots the best solutions in order of positional sensitivity of the lens. These are stored in BestSolutions.mat.

Analysis & Data

Attachment 1: Solutions.pdf
Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf
  1890   Fri Jan 8 17:00:26 2021 AnchalDailyProgressOpticsIncluded lens made by cavity input mirror and distrotion due to crystal

Quote:

Errors in the previous analysis

  • Previously, we wrongly assumed. that the reflected light from the cavity would be as if a reflection is happening from a flat mirror. It actually follows the same paths as incident light in the reflection path.
  • Lens were restricted to non-overlapping regions but that meant that solutions where lens are close to each other can only happen near the boundaries of these regions. Removing this condition widens the search for a good solution.
  • We collimated the beam to near 0.5 mm radius with a 229.1 mm focal length lens at 67" from laser head and put faraday isolator in front. So now the problem only remained to match the mode after this point to the cavity mode.

Goals and restrictions:

  • Use the fewest lenses as possible after having used a fixed lens at 67" point before the faraday isolator.
  • Choose a lens from a list of focal lengths available in west bridge labs.
  • Find the best overlap with the target beam of 18 um at the cavity waist with the most sensitivity with respect to lens positions.
  • The lens should not be closer than 1.5 inches from each other or to the EOM mount or cavity edge.
  • The beam widths should not exceed 4mm in diameter anywhere to ensure small areas of lenses are used.

Analysis & Results


Analysis & Data

 

Attachment 1: Solutions.pdf
Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf
  1889   Thu Jan 7 17:09:16 2021 Paco & AnchalDailyProgressOpticsMode matching OPO

Fresh attempt at mode matching. For this,

  1. Installed the oven, plugged it to the controller and went to the nominal temperature setpoint (40 C) to match the expected path length inside the NL crystal
  2. Placed the output coupler (roc = 15 mm) and roughly align so that the retroreflection is overlapped with the input beam.
  3. Set up a PD (Newfocus 2001) and scope, operating the laser at relatively low power (current ~ 840 mA), and optimize the FI rejected power.
  4. The output coupler is mounted on a three axis mirror mount (Polaris, hoping to get low drift) such that we have some knobs to tune the mode matching initially.

After a couple of iterations moving the mirror X,Y and then scanning all knobs (X,Y, and XY) to effectively translate along Z, the optimized FI rejection is ~(2.15 mW /2.95 mW) 75% of the input beam power. Looking closely at the backreflection from the output coupler, I can clearly see multiple scattered spots, which could definitely account for the defficiency. The most likely culprit is the crystal itself, which is mounted between brass and glass surfaces with no respect for anti-reflection measures. The waist is small enough that no clipping should be happening, so it looks like the NL crystal placement may have to be revisited. Other than that, this procedure should be fine.

  1888   Thu Jan 7 09:44:52 2021 PacoDailyProgressOpticsShaping the OPO cavity mode

Summary of solution number 2 (from previous post). 

After installing the lenses, mirrors and some minor alignment, took the beam profile around the expected minimum waist position (~102" from laser head). The beam profile is astigmatic as can be seen from the plot below (red / blue = x / y), so the mode matching will be suboptimal from the start. 

Taking the geometric mean of the waists (w = sqrt(wx * wy)) we represent our nominal mode and find a min waist of 36.8 um (shaded region in the plot).

The OPO cavity model targets a min waist of 35.5 um (for an optimal Boyd--Klein parameter of ~2.7), but solutions exist with slightly shorter cavities and slightly larger waists which would only compromise the optimal Boyd--Klein parameter to ~2.55 for the sake of better mode matching. I think this is a good place to move out of calculation-land and see how well we can make the cavity work in reality.

Attachment 1: profile_20210106.jpg
profile_20210106.jpg
  1887   Wed Jan 6 09:34:34 2021 AnchalDailyProgressOpticsCorrected analysis and found better solutions.

Errors in the previous analysis

  • Previously, we wrongly assumed. that the reflected light from the cavity would be as if a reflection is happening from a flat mirror. It actually follows the same paths as incident light in the reflection path.
  • Lens were restricted to non-overlapping regions but that meant that solutions where lens are close to each other can only happen near the boundaries of these regions. Removing this condition widens the search for a good solution.
  • We collimated the beam to near 0.5 mm radius with a 229.1 mm focal length lens at 67" from laser head and put faraday isolator in front. So now the problem only remained to match the mode after this point to the cavity mode.

Goals and restrictions:

  • Use the fewest lenses as possible after having used a fixed lens at 67" point before the faraday isolator.
  • Choose a lens from a list of focal lengths available in west bridge labs.
  • Find the best overlap with the target beam of 18 um at the cavity waist with the most sensitivity with respect to lens positions.
  • The lens should not be closer than 1.5 inches from each other or to the EOM mount or cavity edge.
  • The beam widths should not exceed 4mm in diameter anywhere to ensure small areas of lenses are used.

Analysis & Results


Analysis & Data

Attachment 1: Solutions.pdf
Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf
  1886   Thu Dec 31 16:41:59 2020 AnchalDailyProgressOpticsMode matching solution for Cavity

Goals and restrictions:

  • Use the fewest lenses as possible.
  • The beam widths in both onward and reflection direction should be such that there is a 5-inch space somewhere where we can put in the faraday isolator which has an aperture size of 3 mm and intensity limit of 500 W/cm^2.
  • The lens should not be closer than 1.5 inches from each other or to the EOM mount or cavity edge.
  • Choose a lens from a list of focal lengths available in west bridge labs.
  • Find the best overlap with the target beam of 18 um at the cavity waist with the most sensitivity with respect to lens positions.

Analysis & Results


Analysis and Data


Wed Jan 6 10:00:35 2021: This analysis was wrong. See SUS_Lab/1887.

Attachment 1: Solutions.pdf
Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf Solutions.pdf
  1885   Wed Dec 30 09:57:56 2020 PacoDailyProgressOpticsDOPO crystal oven

Assembled first DOPO oven with the crystal. The components (shown below) are:

  • Oven clip
  • Oven
  • ITO crystal spacer
  • PPKTP crystal

The NL crystal sits in the (brass?) clip directly, with the ITO (dummy) crystal pressing it uniformly down. There are no placement references to align the crystal with the oven axis, so this was done very carefully by hand. Once this is roughly straight, the copper arms are fastened in place tight enough to hold everything in place but without excess strain on the NL crystal. The assembly (shown below) is then mounted enclosed in the oven. I put some kapton in place to shield from dust until operation.

Attachment 1: ppktp2.jpg
ppktp2.jpg
Attachment 2: ppktp1.jpg
ppktp1.jpg
  1884   Mon Dec 28 15:51:51 2020 AnchalDailyProgressOpticsMode matching solution for Cavity

Goals and restrictions:

  • Use the fewest lenses as possible.
  • The beam widths in both onward and reflection direction should be such that there is a 5-inch space somewhere where we can put in the faraday isolator which has an aperture size of 3 mm and intensity limit of 500 W/cm2.
  • The lens should not be closer than 1.5 inches from each other or to the EOM mount or cavity edge.
  • Choose a lens from a list of focal lengths available in west bridge labs.
  • Find the best overlap with the target beam of 18 um at the cavity waist with the most sensitivity with respect to lens positions.

Analysis & Results

  • CavityLens.m is run to try all possible lens combinations for 2-lens or 3-lens solutions using ../20201222_BeamProfileNeatEOM/SeedBeam.mat as the seed beam.
  • Then save solutions with more than 70% overlap in CavityModeMatchingSolutions.mat.
  • findBestSolutions.m increases the overlap threshold to 0.9, calculates reflected beam profile for the sideband reflection from the cavity (blue curves in the figures), and discards solution which does not have a 5-inch long area where we can place a faraday isolator with aperture of 3 mm.
  • Black lines show the region where Thorlabs IO-3-1064-HP can be placed safely without clipping or exceeding the intensity limit with 1W power.
  • All solutions in order of sensitivity are plotted here with details of lens choice and positions. In total 7 solutions were found which are stored in BestSolutions.mat.

Analysis & Data


Wed Jan 6 10:00:35 2021: This analysis was wrong. See SUS_Lab/1887.

Attachment 1: Solutions1-7merged.pdf
Solutions1-7merged.pdf Solutions1-7merged.pdf Solutions1-7merged.pdf Solutions1-7merged.pdf Solutions1-7merged.pdf Solutions1-7merged.pdf Solutions1-7merged.pdf
  1883   Wed Dec 23 15:43:01 2020 PacoDailyProgress OPO cavity assembly

Laseroptik optics (4x pairs of cavity mirrors) arrived earlier this week, so I began assembling the input mirror with Noliac (NAC2124) PZT. The (15 mm OD) pzt will sit between a 1" post spacer and the mirror. I applied a thin layer of BT-120-50 (bondatherm) adhesive, which I found in EE shop. From what I gather this adhesive doesn't have softeners (almost doesn't smell) and is a good electrical insulator. The PZT + spacer is sitting below a metallic weight block on the left corner of the table (by the electronics test bench corner), and should finish hardening in a little over 24 hours at room temperature. The PZT was labeled 520 nF (spec. 510 nF).

  1882   Tue Dec 22 15:54:03 2020 AnchalDailyProgressOpticsTook beam profile of near EOM area
  • After installing a 400mm focal length plano-convex lens at 24" from the laser head at (20, 12), we found that higher-order modes are present in the beam.
  • We installed an iris at 34" from laser head at (17, 5)
  • Configuration file is BeamProfileConfiguration2um.ojf.
  • Used fitBeamWidth function of ala mode to fit X and Y beams separately and then their geometric mean.
  • We'll use the geometric mean as the seed profile for future calculations.
  • Found a beam waist of 306 um at 58" from the laser head.
  • Installing the EOM between 49" and 52" from the laser head where the beam waist is between 1 mm and 740 um.

Data

 

Attachment 1: BeamProfileNearEOMAvg.pdf
BeamProfileNearEOMAvg.pdf
  1881   Mon Dec 21 16:35:14 2020 AnchalDailyProgressOpticsTook beam profile of laser right off the head
  • This is a repetition of SUS/1864.
  • Used Data Ray Beam'R2-DD.
  • Took 50 averages and recorded beam "diameters" at 10 different points after the laser head.
  • Configuration file is BeamProfileConfiguration2um.ojf.
  • Used fitBeamWidth function of ala mode to fit X and Y beams separately and then their geometric mean.
  • We'll use the geometric mean as the seed profile for future calculations.

Data

Attachment 1: LaserHeadBeamProfileAvg.pdf
LaserHeadBeamProfileAvg.pdf
Attachment 2: LaserHeadBeamProfileX.pdf
LaserHeadBeamProfileX.pdf
Attachment 3: LaserHeadBeamProfileY.pdf
LaserHeadBeamProfileY.pdf
  1880   Thu Dec 17 12:01:54 2020 PacoLab InfrastructureOpticscrystal ovens, clips and controllers

Covesion order arrived, containing 2x

  • Crystal oven (20 mm long) (below)
  • Clips (for mounting crystals) (below)
  • Blank crystal (to press on the ppktp crystal) (below)
  • OC2 oven controller
  • Controller cable and power cable
Attachment 1: covesion_oven_clip_blank.jpg
covesion_oven_clip_blank.jpg
  1879   Thu Dec 17 11:59:14 2020 PacoMiscEquipmentLoanbeam profiler

Borrow both beam profilers and laptops from WB 264A.

  1878   Wed Dec 16 22:38:55 2020 KojiMiscEquipmentLoanTook delay line box, compressed nitrogen cylinder and lens

Photos, please, because we don't allow a free-rolling cylinder in a lab.

ELOG V3.1.3-